What is the answer to -6+5x=8x-9

Answer by checkley71(8403)

-6+5X=8X-9
5X-8X=-9+6
-3X=-3
X=-3/-3
X=1 ANSWER.
PROOF
-6+5*1=8*1-9
-6+6=9-9
0=0